# Safe and Economical UAV Trajectory Planning in Low-Altitude Airspace: A Hybrid DRL-LLM Algorithm with Compliance Awareness

### v2

@@ 3.1 System Model
Data Collection UAV (DCU). The DCU is responsible for the data collection task in the low-altitude economy and flies at a fixed altitude $H$. It takes off from $\mathbb{T}\mathbb{A}$, then conducts data collection in $\mathbb{A}$, and eventually reaches $\mathbb{L}\mathbb{A}$. When conducting data collection, the DCU is not allowed to fly into NFZs. Furthermore, when the DCU flies into residential zones (RZs), to ensure noise control, its flight speed must be less than $v_{\text{limit}}$} [28]. The DCU must also ensure that it does not collide with building zones (BZs) and other UAVs (OUs) in the low-altitude economic networking. Therefore, we assume that the DCU is equipped with a sensing device with a perception radius of $PR$. Through this device, the DCU can detect the positions and flight directions of OUs and thereby adjust its flight trajectory to avoid collisions.
@@
Ground Equipment (GE). GE refers to devices that generate data in the low-altitude economy, such as intelligent traffic devices, power infrastructure equipment, and environmental monitoring devices [37]. These data are collected regularly by the DCU.
@@ 3.2 Models Related to UAV
This section presents models about how UAV moves, transmits data, and consumes energy in our system. The high mobility, dynamic topology, and fast-fading channels cause Doppler frequency shift [41], which induces intercarrier interference and raises the bit error rate, thereby degrading DCU data collection performance. Therefore, we use the “hover-and-fly” scheme for the DCU [32].
@@ 3.2.1 Mobility Model
The movement model is used to describe how the position of the DCU changes at times $t$ and $t+\Delta t$, that is, how the DCU moves. Given the DCU’s position ${L_{u,t}}=({x_{u,t}},{y_{u,t}},{z_{u,t}})$ and the velocity ${v_{u}}(t)=({v_{u}}{(t)_{x}},{v_{u}}{(t)_{y}},{v_{u}}{(t)_{z}})$ at time $t$, the DCU’s position at time $t+\Delta t$ is as follows:
${L_{u,t+\Delta t}}=$ (1)
$({x_{u,t}}+{v_{u}}{(t)_{x}}{t_{fly}},\ {y_{u,t}}+{v_{u}}{(t)_{y}}{t_{fly}},\ {z_{u,t}}+{v_{u}}{(t)_{z}}{t_{fly}}).$
Both DCU and GE are equipped with a single antenna [43]. The entire data collection period is divided into equal-length time slots ($\{{t_{i}}\}_{i=1}^{T/\Delta t}$). In a given time slot $\Delta t$, the process is split into two phases: a flight phase, where the DCU searches for the optimal communication position, and a data collection phase, where it hovers at that position to collect data. Therefore, given ${t_{fly}}=\Delta t/2$, ${z_{u,t}}=H$, ${v_{u}}{(t)_{z}}=0$, the distance between $L_{u,t+\Delta t}$} and $L_{u,t}$} should satisfy the following constraint:
$\left\|{L_{u,t+\Delta t}}-{L_{u,t}}\right\|\leq{v_{\max}}{t_{fly}}.$ (2)
If the DCU flies into $\mathbb{R}_{RZ}$}, the distance between $L_{u,t+\Delta t}$} and $L_{u,t}$} should satisfy the following constraint:
$\left\|{L_{u,t+\Delta t}}-{L_{u,t}}\right\|\leq{v_{{\rm{limit}}}}{t_{fly}}.$ (3)
Note that the mobility model of OUs is the same as that of DCU.
@@ 3.2.2 Data Transmission Model
The data transmission model describes how much the data collection volume of the DCU changes at times $t$ and $t+\Delta t$, that is, how much data the DCU can transmit within time $\Delta t$ ($T{D_{u,G{E_{i}}}}(t)$). Before giving how to compute $T{D_{\Delta t}}$, we first present the effective information rate of the DCU in the time slot $\Delta t$. The DCU collects data from a single randomly selected GE within its communication range in each time slot by using the standard time division multiple access [38]. When DCU collects data, it is in a hovering state without moving. The DCU begins to collect data at time $t+{t_{fly}}$ ($t_{fly}=\Delta t/2$) and lasts for ${t_{col}}=\Delta t-{t_{fly}}$. As a result, assuming the ground device that is currently collecting data is $GE_{i}$, the effective information rate during the time slot $\Delta t$ is as follows:
${B_{u,GE_{i}}}(t)=\log_{2}\left(1+SNR_{u,GE_{i}}(t)\right),$ (4)
where $SNR_{u,GE_{i}}(t)=TP_{GE_{i}}\cdot G_{u,GE_{i}}(t)/\mathcal{N}$ is the signal-to-noise ratio in the time slot $\Delta t$ beginning at time $t$, $TP_{GE_{i}}$ is the transmit power of $GE_{i}$, $G_{u,GE_{i}}(t)$ is the channel power gain from the DCU to $GE_{i}$ in the time slot $\Delta t$ beginning at time $t$, and $\mathcal{N}$ is the noise power.
Given the flight altitude of the DCU, the communication link with $GE_{i}$ is assumed to follow a line-of-sight channel model [51]. The free-space path-loss model is used to compute $G_{u,GE_{i}}(t)$ as follows [44]:
$G_{u,GE_{i}}(t)=$ (5)
@@
where $\alpha$ is the channel power gain at $\sqrt{\left\|(x_{u,t+t_{fly}},y_{u,t+t_{fly}},H)-(x_{GE,i},y_{GE,i},0)\right\|^{2}}=1$ m distance.
Note that, if the DCU wants to transmit data, $SNR_{u,GE_{i}}(t)$ must exceed a certain threshold $\tau$ [45]. Therefore, according to Equations (4) and (5), $TD_{u,GE_{i}}(t)$ can be computed as follows:
@@ 3.2.3 Energy Consumption Model
The energy consumption model is used to describe how much the remaining energy of the the DCU changes at times $t$ and $t+\Delta t$, that is, how much energy the DCU consumes within time $\Delta t$ ($E(t)$). The energy consumption of the DCU within $\Delta t$ beginning at time $t$ can be divided into two parts: $e_{\text{fly}}(t)$ and $e_{\text{hover}}(t)$. For $e_{\text{fly}}(t)$, it can be computed as follows [47]:
@@
$\Big{(}\frac{\lambda}{8}\cdot\rho\eta A_{r}\Im^{3}r_{\text{rotor}}^{3}\left(1+\frac{3v_{u}(t)^{2}}{U_{\text{tip}}^{2}}\right)+\frac{1}{2}d_{0}\rho sA_{r}v_{u}(t)^{3}\Big{)}\cdot t_{\text{fly}},$
where $\lambda$ denotes the profile drag coefficient, $A_{r}$ the rotor disc area, $\Im$ the blade angular velocity, $r_{\text{rotor}}$ the rotor radius, $d_{0}$ denotes the fuselage drag coefficient, $\rho$ the density of air, $\eta$ the rotor solidity factor, $U_{\text{tip}}$ the tip velocity of the rotor blade, $\lambda$ the average induced velocity during hovering, and $v_{u}$ the velocity of DCU.
@@
$\left((1+\Re)w^{3/2}\sqrt{1+\frac{v_{u}(t)^{4}}{4v_{0}^{4}}}-\frac{v_{u}(t)^{2}}{2v_{0}}\middle/\sqrt{2\rho A_{r}}\right)\cdot t_{\text{hover}},$
@@ 4.1 Requirement Analysis
R1) Limited Mobility. The moving distance of DCU during each time slot should satisfy the Equation  (2).
@@
R2) Robust Obstacle Avoidance. Since data collection in the low-altitude economy is concentrated in urban areas and there is airspace reuse in low-altitude areas [36], when the DCU conducts data collection, it not only needs to avoid ${\mathbb{R}}_{BZ}$}, ${\mathbb{R}}_{NFZ}$}, but also OUs.
R3) Compliance. Unlike general data collection tasks, DCU operating in urban low-altitude environments must comply with strict regulations. They must reduce speed near to minimize noise and meet regulatory standards.
R4) Economy. Within the low-altitude economy, data collection tasks must jointly consider the maximization of data acquisition and the minimization of energy consumption to maximize the data collection task benefits. Consequently, the DCU is required not only to collect sufficient data but also to complete the task along an energy-efficient trajectory that minimizes energy costs.
@@ 4.2 Problem Formulation
$\textbf{OP:}\quad\arg\max\sum\nolimits_{t=0}^{T}TD_{u,GE_{i}}(t),$ (9)
@@
$GE_{i}=\arg\max\left(\{B_{u,GE_{i}}(t)\}_{i=1}^{N_{GE}}\wedge SNR_{u,GE_{i}}(t)\geq\tau\right),$ (9a)
$t+1=t+\Delta t,\quad\forall t,$ (9b)
$L_{u,t+1}=(x_{u,t}+v_{u}(t)_{x}\cdot t_{\text{fly}},\ y_{u,t}+v_{u}(t)_{y}\cdot t_{\text{fly}},\ z_{u,t}),$ (9c)
$v_{u}(t)\leq v_{\max},\quad\forall t,$ (9d)
$v_{u}(t)\leq v_{\text{limit}},\quad\text{if }L_{u,t}\in\{\mathbb{R}_{RZ,m}\}_{m=1}^{N_{RZ}},$ (9e)
$\begin{aligned} &(L_{u,t},L_{u^{\prime},t})\notin\{\mathbb{R}_{BZ,n}\}_{n=1}^{N_{BZ}}\\ &\wedge\ (L_{u,t},L_{u^{\prime},t})\notin\{\mathbb{R}_{NFZ,l}\}_{l=1}^{N_{NFZ}},\quad\forall t,\end{aligned}$ (9g)
$\left\|L_{u,t}-L_{u^{\prime}_{j},t}\right\|>0,\quad\forall t,\ j\in[1,N_{U^{\prime}}],$ (9h)
$\sum\nolimits_{t=0}^{T}(e_{\text{fly}}(t)+e_{\text{hover}}(t))\leq E_{\text{limit}}.$ (9i)
The objective function OP maximizes the amount of data collected over the given time slots. Constraint (9a) selects a GE to collect data because the DCU follows a maximum signal strength strategy to connect to the GE with the highest received signal power [51], while (9b) presents the relationship between $t$ and $t+1$ according to the data transmission model. Constraint (9c) governs the position updates of the DCU according to the mobility model. Constraint (9d) enforces the maximum speed limit, and Constraint (9e) enforces the maximum speed limit in $\mathbb{R}_{RZ}$ to satisfy R1 and R3. Constraints (9f), (9g), and (9h) ensure that the DCU operates within $\mathbb{A}$, avoids $\mathbb{R}_{NFZ}$, and prevents collisions with $\mathbb{R}_{BZ}$ or OUs to satisfy R2 and R3. Constraint (9i) ensures the maximum energy consumption of the DCU in a data collection task to satisfy R4.

### v3

@@ 3.1 System Model
Data Collection UAV (DCU). The DCU is responsible for the data collection in the low-altitude economy and flies at a fixed altitude $H$. It takes off from $\mathbb{T}\mathbb{A}$, then conducts data collection in $\mathbb{A}$, and eventually reaches $\mathbb{L}\mathbb{A}$. When conducting data collection, the DCU is not allowed to fly into NFZs. Furthermore, when the DCU flies into residential zones (RZs), to ensure noise control, its flight speed must be less than $v_{\text{limit}}$} [25]. The DCU must also ensure that it does not collide with building zones (BZs) and other UAVs (OUs) in the low-altitude economic networking. Therefore, we assume that the DCU is equipped with a sensing device with a perception radius of $PR$. Through this device, the DCU can detect the positions and flight directions of OUs and thereby adjust its flight trajectory to avoid collisions.
@@
Ground Equipment (GE). GE refers to devices that generate data in the low-altitude economy, such as intelligent traffic devices, power infrastructure equipment, and environmental monitoring devices [33]. These data are collected regularly by the DCU. $N_{\mathrm{GE}}$ GEs are placed independently and uniformly at random over $\mathbb{A}$, while maintaining appropriate clearance from obstacle boundaries and zones (NFZ/BZ/RZ). Such random deployment is widely adopted in UAV-enabled data collection trajectory planning scenarios, as it captures diverse and unpredictable real-world layouts without imposing artificial structural patterns.
@@ 3.2 Models Related to UAV
This section presents models about how UAV moves, transmits data, and consumes energy in our system.
@@ 3.2.1 Mobility Model
The movement model is used to describe how the position of the DCU changes. The high velocity and dynamic topology of UAVs can lead to Doppler frequency shift (DFS) and fast-fading effects, which significantly degrade communication quality during data collection [37]. To mitigate these effects and maintain stable links with mobile devices, DCU adopts a hover-and-fly strategy [38]. In this scheme, the DCU first travels within the designated area $\mathbb{A}$ to locate an optimal communication position and then hovers at that point to collect data. This hovering mechanism effectively stabilizes the communication channel by reducing DFS-induced distortion and intercarrier interference.
To model this process in a tractable manner, the entire trajectory planning period is divided into equal time slots of length $\Delta t$, where $t\in[0,T]$ denotes the slot index and $T$ is the total number of time slots determined by mission duration. Each time slot is equally divided into a flight phase and a data collection phase. During the flight phase, the DCU moves within $\mathbb{A}$ to search for a better position, and during the data collection phase, it hovers to communicate with mobile devices. In practice, there may be cases where no devices are within the sensing range during data collection, causing idle waiting and unnecessary energy consumption.
To minimize this energy waste and maintain modeling simplicity, the duration of both phases is set to 1 s, yielding a total slot length of $\Delta t=2$ s. This fine-grained time division allows the DCU to frequently adjust its position while ensuring stable communication in each slot. The DCU is equipped with a single antenna and moves within the area $\mathbb{A}$. The DCU flies at a fixed altitude $H$ with a maximum speed $v_{\max}$ and can communicate with at most one device per slot.
Given the DCU’s position ${L_{\mathrm{DCU},t}}=({x_{\mathrm{DCU},t}},{y_{\mathrm{DCU},t}},H$ and the velocity ${v_{\mathrm{DCU}}}(t)=(v_{x}(t),v_{y}(t),0)$ at the beginning of time slot $t$, the DCU’s position at time $t+1$ is as follows:
${L_{\mathrm{DCU},t+1}}=$ (1)
$({x_{\mathrm{DCU},t}}+v_{x}(t){t_{\mathrm{fly}}},\ {y_{\mathrm{DCU},t}}+v_{y}(t){t_{\mathrm{fly}}},\ H),$
where ${t_{\mathrm{fly}}}=\Delta t/2$ is the length of the flight phase. Given the maximum flight speed of the DCU, the distance between $L_{\mathrm{DCU},t+1}$} and $L_{\mathrm{DCU},t}$} should satisfy the following constraint:
$\left\|{L_{\mathrm{DCU},t+1}}-{L_{\mathrm{DCU},t}}\right\|\leq{v_{\max}}{t_{\mathrm{fly}}}.$ (2)
If the DCU flies into $\mathbb{R}_{\mathrm{RZ}}$}, the distance between $L_{\mathrm{DCU},t+1}$} and $L_{\mathrm{DCU},t}$} should satisfy the following constraint:
$\left\|{L_{\mathrm{DCU},t+1}}-{L_{\mathrm{DCU},t}}\right\|\leq{v_{{\mathrm{limit}}}}{t_{\mathrm{fly}}},$ (3)
where $v_{\mathrm{limit}}$ is the speed limit in ${\mathbb{R}}_{\mathrm{RZ}}$} for the DCU.
The OUs follow the same mobility model as the DCU described in Eq. (1). However, unlike the DCU whose trajectory is optimized by the proposed algorithm, the trajectories of all OUs are randomly generated prior to simulation. Each OU’s velocity $v_{\mathrm{OU}}(t)$ is sampled from a uniform distribution within $0,v_{\max}]$, and its movement direction is randomly initialized and updated according to the same motion equations as the DCU. During trajectory generation, each OU’s path is constrained to ensure that it does not intersect with any BZs or NFZs, thereby maintaining physically valid and collision-free motion throughout the environment. This stochastic trajectory generation emulates the unpredictable movement of other UAVs or aerial entities in low-altitude airspace, introducing dynamic uncertainty for the DCU’s trajectory planning.
@@ 3.2.2 Data Transmission Model
The data transmission model describes how much the data collection volume of the DCU collects in the time slot $t$. Before giving how to compute it, we first present the effective information rate of the DCU in the time slot $t$. The DCU collects data from a single selected GE within its communication range in each time slot by using the standard time division multiple access [34]. When DCU collects data, it is in a hovering state without moving. The DCU begins to collect data at time $t+{t_{\mathrm{fly}}}$ ($t_{\mathrm{fly}}=\Delta t/2$) and lasts for ${t_{\mathrm{hover}}}=\Delta t-{t_{\mathrm{fly}}}$. As a result, assuming the GE that is currently collecting data is $\mathrm{GE}_{i}$, the effective information rate during the time slot $t$ is as follows:
${B_{\mathrm{GE}_{i}}}(t)=\log_{2}\left(1+{\mathrm{SNR}}_{\mathrm{GE}_{i}}(t)\right),$ (4)
where ${\mathrm{SNR}}_{\mathrm{GE}_{i}}(t)=TP_{\mathrm{GE}_{i}}\cdot G_{\mathrm{GE}_{i}}(t)/\mathcal{N}$ is the signal-to-noise ratio at the beginning of time slot $t$, $TP_{\mathrm{GE}_{i}}$ is the transmit power of $\mathrm{GE}_{i}$, $G_{\mathrm{GE}_{i}}(t)$ is the channel power gain from the DCU to $\mathrm{GE}_{i}$ at the beginning of time slot $t$, and $\mathcal{N}$ is the noise power.
Given the fixed flight altitude of the DCU $H$, the communication link with $\mathrm{GE}_{i}$ is assumed to follow a line-of-sight (LoS) air-to-ground (A2G) channel model [46]. This assumption aligns well with the characteristics of UAV-assisted data collection, where the DCU actively approaches each GE and performs data acquisition at short distances during hovering. Such close-range outdoor A2G communication typically presents dominant LoS components, making the LoS model a widely adopted and effective abstraction in UAV data collection trajectory planning scenarios. The free-space path-loss model is used to compute $G_{\mathrm{GE}_{i}}(t)$ as follows [39]:
$G_{\mathrm{GE}_{i}}(t)=$ (5)
@@
where $\alpha$ is the channel power gain at $\sqrt{\left\|(x_{\mathrm{DCU},i+t_{\mathrm{fly}}},y_{\mathrm{DCU},i+t_{\mathrm{fly}}},H)-(x_{{\mathrm{GE}}_{i}},y_{{\mathrm{GE}}_{i}},0)\right\|^{2}}=1$ m distance.
Note that, if the DCU wants to transmit data, ${\mathrm{SNR}}_{\mathrm{GE}_{i}}(t)$ must exceed a certain threshold $\tau$ [40]. Therefore, according to Eqs. (4) and (5), the data collection volume of the DCU collects in the time slot $t$ $TD_{\mathrm{GE}_{i}}(t)$ can be computed as follows:
@@
where $DV_{\mathrm{GE}_{i},t}$ is the data volume of the $i$-th GE at the beginning of time slot $t$.
@@ 3.2.3 Energy Consumption Model
The energy consumption model is used to describe how much the remaining energy of the the DCU changes in the time slot $t$. The energy consumption of the DCU in the time slot $t$ can be divided into two parts: the energy consumed by flying (denoded as $e_{\text{fly}}(t)$) and the energy consumed by hovering (denoted by $e_{\text{hover}}(t)$). For $e_{\text{fly}}(t)$, it can be computed as follows [42]:
@@
$\Big(\frac{\lambda}{8}\rho\eta A_{r}\Im^{3}r_{\text{rotor}}^{3}\left(1+\frac{3v_{\mathrm{DCU}}(t)^{2}}{U_{\text{tip}}^{2}}\right)+\frac{1}{2}d_{0}\rho sA_{r}v_{\mathrm{DCU}}(t)^{3}\Big)t_{\text{fly}},$
where $\lambda$ denotes the profile drag coefficient, $A_{r}$ the rotor disc area, $\Im$ the blade angular velocity, $r_{\text{rotor}}$ the rotor radius, $d_{0}$ denotes the fuselage drag coefficient, $\rho$ the density of air, $\eta$ the rotor solidity factor, $U_{\text{tip}}$ the tip velocity of the rotor blade, and $v_{\mathrm{DCU}}$ the velocity of DCU.
@@
$\left((1+\Re)w^{3/2}\sqrt{1+\frac{v_{\mathrm{DCU}}(t)^{4}}{4v_{0}^{4}}}-\frac{v_{\mathrm{DCU}}(t)^{2}/2v_{0}}{\sqrt{2\rho A_{r}}}\right)t_{\text{hover}},$
@@ 4.1 Requirement Analysis
R1) Limited Mobility. The moving distance of DCU during each time slot should satisfy the Eq. (2).
@@
R2) Robust Obstacle Avoidance. Since data collection in the low-altitude economy is concentrated in urban areas and there is airspace reuse in low-altitude areas [32], when the DCU conducts data collection, it not only needs to avoid ${\mathbb{R}}_{\mathrm{BZ}}$}, ${\mathbb{R}}_{\mathrm{NFZ}}$}, but also OUs.
R3) Compliance. Unlike general trajectory planning tasks, DCU operating in urban low-altitude environments must comply with strict regulations. They must reduce speed near ${\mathbb{R}}_{\mathrm{RZ}}$} to minimize noise and meet regulatory standards.
R4) Economy. Within the low-altitude economy, trajectory planning tasks must jointly consider the maximization of data acquisition and the minimization of energy consumption to maximize the trajectory planning task benefits. Consequently, the DCU is required not only to collect sufficient data but also to complete the task along an energy-efficient trajectory that minimizes energy costs.
@@ 4.2 Problem Formulation
$\textbf{OP:}\quad\underset{v_{\mathrm{DCU}}(t)}{\arg\max}\;\sum\nolimits_{t=1}^{T}TD_{\mathrm{GE}_{i}}(t)$ (9)
@@
$\mathrm{GE}_{i}=\arg\max\left(\{B_{\mathrm{GE}_{i}}(t)\}_{i=1}^{N_{\mathrm{GE}}}\wedge({\mathrm{SNR}}_{\mathrm{GE}_{i}}(t)\geq\tau)\right),$ (9a)
$L_{\mathrm{DCU},t+1}$
$=(x_{\mathrm{DCU},t}+v_{x}(t)t_{\text{fly}},\ y_{\mathrm{DCU},t}+v_{y}(t)t_{\text{fly}},\ H),$ (9b)
$v_{\mathrm{DCU}}(t)=(v_{x}(t),v_{y}(t),0)\leq v_{\max},\quad\forall t,$ (9c)
$v_{\mathrm{DCU}}(t)\leq v_{\text{limit}},\quad\text{if }L_{\mathrm{DCU},t}\in\{\mathbb{R}_{\mathrm{RZ},\mathit{m}}\}_{m=1}^{N_{\mathrm{RZ}}},$ (9d)
$(L_{\mathrm{DCU},t},L_{\mathrm{OU},t})\in\mathbb{A},\quad\forall t,$ (9e)
$(L_{\mathrm{DCU},t},L_{\mathrm{OU},t})\notin\{\mathbb{R}_{\mathrm{NFZ},\mathit{l}}\}_{l=1}^{N_{\mathrm{NFZ}}},\quad\forall t,$ (9f)
$(L_{\mathrm{DCU},t},L_{\mathrm{OU},t})\notin\{\mathbb{R}_{\mathrm{BZ},\mathit{n}}\}_{n=1}^{N_{\mathrm{BZ}}},\quad\forall t,$ (9g)
$\left\|L_{\mathrm{DCU},t}-L_{\mathrm{OU}_{j},t}\right\|>0,\quad\forall t,\ j\in[1,N_{\mathrm{OU}}],$ (9h)
$(L_{\mathrm{DCU},1}=\mathrm{SL}\in\mathbb{TA})\wedge(L_{\mathrm{DCU},T}=\mathrm{EL}\in\mathbb{LA}),$ (9i)
$\sum\nolimits_{t=1}^{T}(e_{\text{fly}}(t)+e_{\text{hover}}(t))\leq E_{\text{limit}},\quad\forall t.$ (9j)
The objective function OP maximizes the amount of data collected over the given time slots. Constraint (9a) selects a GE to collect data because the DCU follows a maximum signal strength policy to connect to the GE with the highest received signal power [46]. Note that constraint (9a) defines the feasible region for data collection, where a GE can be served only when the DCU is within its communication range, instead of enforcing a distance-priority rule. Once the DCU enters the feasible range of a GE, data collection occurs automatically. When multiple GEs have the same effective rate $B_{\mathrm{GE}_{i}}(t)$, a randomly GE will be collected data by the DCU. Constraint (9b) governs the position updates of the DCU according to the mobility model. Constraint (9c) enforces the maximum speed limit, and Constraint (9d) enforces the maximum speed limit in $\mathbb{R}_{\mathrm{RZ}}$ to satisfy R1 and R3. Constraints (9e), (9f), (9g), and (9h) ensure that the DCU operates within $\mathbb{A}$, avoids $\mathbb{R}_{\mathrm{NFZ}}$, and prevents collisions with $\mathbb{R}_{\mathrm{BZ}}$ or OUs to satisfy R2 and R3. constraint (9i) ensures DCU takes off from $\mathbb{TA}$ and lands off in $\mathbb{LA}$, while constraint (9j) ensures the maximum energy consumption of the DCU in a trajectory planning task to satisfy R4.